OpenStudy (anonymous):

Please Help!! Will Medal!! The location of a point on a coordinate plane is given by an ordered pair (m, n), where m and n are integers such that -5 d m d 5 and -5 d n d 5. How many different points fit that description? A. 121 B. 100 C. 22 D. 1331

OpenStudy (aric200):

:I honestly im clueless haha

OpenStudy (anonymous):

Tis alright, I just need to get all this work over with.

OpenStudy (aric200):

Same bro,this is my break from doing 20 essays due tomorrow...only 12 more left :'D

OpenStudy (anonymous):

20 Essays??????

OpenStudy (aric200):

My research class is due by 12:a.m tomorrow so yep 12 left

OpenStudy (anonymous):

I have like a day to get in 15 math assignments at 10 q's each....And I haven't even touched the finals.

OpenStudy (anonymous):

I'm a killer writer if you need help

OpenStudy (aric200):

You are in online school I assume then?

OpenStudy (anonymous):

Nah just this one class. Messed me over.

OpenStudy (anonymous):

Have to graduate Monday. So I need all this in by Thursday.

OpenStudy (aric200):

Oh pellet yeah get it done.

OpenStudy (aric200):

Only a sophomore here.

OpenStudy (aric200):

pellet haha

OpenStudy (anonymous):

Lol, take some advice from me. Do not procrastinate.

OpenStudy (aric200):

:I yeah If only i met you about eh 4 months ago and I would not have

OpenStudy (aric200):

I havent touched biology and its 208 assignments due in 4 weeks...Fun times

OpenStudy (aric200):

I am off but good luck with your stuff sorry i could not have been more of help I need to get back to work, hope all goes well I'll see you around.Deuces.

OpenStudy (anonymous):

Fuuuuuck that oh man. Alright dude good luck D:

ganeshie8 (ganeshie8):

Ho many choices are there for \(m\) ? Ho many choices are there for \(n\) ?

OpenStudy (anonymous):

-5 d m d 5 and -5 d n d 5.

ganeshie8 (ganeshie8):

thats not readable :/

ganeshie8 (ganeshie8):

could u take a snapshot of original question and attach ? @moemand

OpenStudy (anonymous):

this one is hard I have no Idea

OpenStudy (anonymous):

):

OpenStudy (anonymous):

Yes @ganeshie8

ganeshie8 (ganeshie8):

Okay there is problem wid syntax in the quesiton, the question actually was saying this : \(-5 \le m \le 5\) and \(-5 \le n \le 5\)

ganeshie8 (ganeshie8):

That means there are \(11\) choices for each of \(m\) and \(n\)

OpenStudy (anonymous):

Yeah I didn't understand this at all, but okay so 11 choices for m and n

ganeshie8 (ganeshie8):

So total possible choices for the ordered pair \((m, n)\) = \(11 \times 11 = 121\)

OpenStudy (anonymous):

Thank you!!!!

ganeshie8 (ganeshie8):

yw, btw the number of choices are 11 cuz the integers between -5 and 5 are 11 : `-5, -4, -3, -2, -1, 0, 1, 2, 3, 4, 5` count'em up